A pendulum bod weighing 500 gm is suspended on a fine wire. what is the tension on the wire when the bod wings at an angle of 60 degree from the vertical?


1
Expert's answer
2021-09-04T07:35:15-0400
"\sum F_y=0:\quad T\cos 60^{\circ}-mg=0"

"T=mg/\cos 60^{\circ}\\
=0.500\:\rm kg*9.8\: N/kg/\cos 60^{\circ}=9.8\: N"